A racquetball thrown from the ground at an angle of 28° and with a speed of 35 m/s lands exactly 1.6 s later on the top of a nearby building. Calculate the horizontal distance it traveled and velocity just before hitting the building.


1
Expert's answer
2021-01-26T08:42:55-0500

The horizontal distance it traveled:


"R=v_xt=v\text{ cos}28°t=49\text{ m}."

Find the velocity just before hitting the building. First, calculate the time it takes to reach the highest point of trajectory:


"\tau=\frac{v_y}{g}=\frac{v\text{ sin}28°}{g}=1.6767\text{ s}."

It means that after 1.6 s, the object had certain vertical component of velocity, which equals


"v_{y1}=v_y-gt=v\text{ sin}28°-gt=0.75\text{ m/s}."

So, the magnitude of velocity right before hitting the building is


"v(t)=\sqrt{v_x^2+v_{y1}^2}{}=30.9\text{ m/s}."